A stock had annual returns of 8 percent, 14 percent, and 2 percent for the past three years. Based on these returns, what is the probability that this stock will return more than 26 percent in any one given year?
Answer #1

Mean= (8%+14%+2%)/3 = 8%

Standard deviation (SD) = {1/(3-1)*[(8%-8%)^2 + (14%-8%)^2 + (2%-8%)^2]}^0.5 = 6%

Stock has to be more than 26% so z-score is (26%-mean)/SD = (26%-8%)/6% = 3

Reading from the z-score table, probability that the return will be less than 26% is 99.87% so probability that it is more than 26% is 100%-99.87% = 0.13%